Let f(x)=2−x and g(x)=x3-2, find:


a.(f∘g)(0)=


b. (g∘f)(0)=


c. (f∘g)(x)=

     

d. (g∘f)(x)=


1
Expert's answer
2021-03-11T06:18:42-0500

a.

"(f\circ g)(x)=2-(x^3-2)=4-x^3"




"(f\circ g)(0)=4-(0)^3=4"



b.

"(g\circ f)(x)=(2-x)^3-2"


"(g\circ f)(0)=(2-0)^3-2=6"

c.

"(f\circ g)(x)=2-(x^3-2)=4-x^3"



g.

"(g\circ f)(x)=(2-x)^3-2"